1. A ball rolls off the edge of a 1.2 m high and strikes the floor at a point 2 m horizontally from the edge of the table. Find the (a) time of its flight (b) its initial horizontal velocity.
1
Expert's answer
2020-11-10T07:03:02-0500

a)


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(1.2)}{9.8}}=0.49\ s"

b)


"v_x=\frac{x}{t}=\frac{2}{0.49}=4.1\frac{m}{s}"